PHYSICS 2140 INTRODUCTORY PHYSICS

FALL 1999 QUIZ SECTION

Quiz Instructor:  Robert Harr  Office:  341 Physics
Phone:  577-2677  E-mail:  harr@physics.wayne.edu
Office Hours:  5–6 TTh, 9:30–10:30 MW or by appointment
Quiz Section:  10:40–11:35 T or Th, 177 Physics

Quiz periods give you a chance to ask questions about course material and problems. I hope that the quiz section can be “student driven”, though I will normally prepare a small amount of material.

Quizzes: There will be several unannounced quizzes during the semester. A quiz will normally be given during the last ten or fifteen minutes of the quiz period. More details about the quiz grades will be given by Dr. Padmanabahn.